OpenStudy (anonymous):

multiply: sq root of 5 * sq root of 2=

OpenStudy (anonymous):

sqrt10

OpenStudy (anonymous):

sqrt{5}sqrt{2} sqrt{10}

OpenStudy (anonymous):

Ha! I beat pk

OpenStudy (anonymous):

Congrats!!!

OpenStudy (anonymous):

said review product rule for radicals??

OpenStudy (anonymous):

sqrt{ab} = sqrt{a}sqrt{b}

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\]

OpenStudy (anonymous):

so why is 10 the wrong answer?

OpenStudy (anonymous):

because you forgot to put the square root symbol above it

OpenStudy (anonymous):

ohhhhh.... I think I know....

OpenStudy (saifoo.khan):

sqrt10.